OpenStudy (anonymous):

A weather balloon filled with helium has a radius of 1m and a mass of 1 kg (linked to ropes and the latex balloon). What is the maximum load of the scientific package this balloon can raise when it floats at an altitude where the helium density is 0.16 kg/m3 and the air density is 1.2kg/m3? (assume that only the balloon displaces air).

OpenStudy (anonymous):

this question involves use of archimedes principle upward buoyant force that keeps ballon afloat is >or equal to weight of the baloon so writing it in terms of expression: density of baloon=Db ; density of air=Da Da*Va*g=Db*Vb*g i equate as i want to find the MAXIMUM loiad on balloon i lose my g Va=volume of air displaced by baloon=volume of balloon(volume of sphere)=4/3*pi*r^3 r=1m now find Vb now i got the volume of balloon (with its contents inside that makes it heavy) Vb=mass/density Db so find mass then subtract 1 kg from it to find weight of load the thinking pattern for the question is: i want the balloon to reemain floating so what should i do(the air does) to keep it afloat next ask whether u can manage putting some expressions for handling the same